At L=4, K=4, the marginal product of labour is 2 and the marginal product of capital is 3. What is the marginal rate f technical substitution?
1
Expert's answer
2015-03-31T10:42:49-0400
At L=4, K=4 MPl = 2, MPk = 3.
Marginal Rate of Technical Substitution (MRTS) is the amount by which the quantity of one input has to be reduced when one extra unit of another input is used, so that output remains constant. MRTS(K, L) = MPk/MPl = 2/3 = 0.67
So, the marginal rate of technical substitution is 0.67.